. Find the square of the following numbers.  iii. 86  
,jg,kymtdjrdtxhyrfjkhyktjuryhrdjyk,y

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

iii. (86)2
=(90-4)2
= (90)2+ (4)2- 2×90×4 =8100+16–720 =8116–720
= 7396



Thanks